Urea is stronger base than ordinary amide?according to lewis theory

Grade:12th pass

3 Answers

Umakant biswal
7 years ago
amide are weak base due to the resonance of the amide group . 
in acetamide , the lone pair nitrogen are not avaliable for bonding , with h ^+ . 
however in urea , the lone pair of second nh2 group avaliable some of the time for bonding with h+ .. 
so, it will be stronger base than acetamide . 
HOPE IT CLEARS 
ALL THE BEST ..
Vikas TU
7 years ago
The reason is:
As the lone pair in Urea is freely available and thus not participates in any Resonance.
But in any ordinary amide the lone pair is conjugated with the resoace and hence hindered to donate elctrons thus weak base as compare to Urea.
jyoti bhatia
7 years ago
 
Urea has two nitrogen having two lone pair of electrons. Acetamide had only one nitrogen and only one li=one pair of electrons. 
Urea is a stronger base because it has lower pKb than acetamide. reason for this is:
 
Amides are weak bases due to resonance of amide group.
R–C(=O)–NH2 ↔ R–C(–O^-)=NH2+
In actamide (R=CH3), lone pair of nitrogen is not available for bonding with H^+. However, in urea (R+NH2), lone pair on second NH2 group is available for bonding with H^+ (hydrogen ions), so it is a stronger base than acetamide.
Acetamide: pKb = 13.5
Urea: pKb = 13.8